Fermi National Laboratory
 

Accelerator Update for August 10, 2007 - August 17, 2007

 

Fermilab Shutdown Report

8/10/07 – 8/17/07

LINAC

Booster

Electrical Work:

  • Cable pull resumed Thursday 08/16 (~ 70% complete)

Water work:

  • LCW water work in the Booster West Gallery
  • Much work in area – walk around
  • Installation of 6” pipe ongoing
  • Work started at the south end of the gallery going north
  • Next week: Continue with the installation
      • 4” and 6” copper pipe are staged in the gallery.
  • CUB VFD testing was performed on Thursday and need to be redone.

Corrector Installation:

  • Number of correctors in the tunnel: 5
    • 4 fully surveyed and vacuum connections complete (L19, L17, L21, & L23)
    • Physically installed corrector at L15
      • Due to a technical problem discovered during the corrector installation at L15, the magnet support plates had to be shipped to machine shop for necessary adjustments
    • Prep L11 for corrector installation today
  • 5 more correctors were delivered to Hi Bay area (BMA013, BMA020, BMA021, BMA023, BMA024)
    • 2 more are needed in order to have all 12 on hand.
  • Next week: Continue with corrector installation

Antiproton Source

Mechanical

  • Leak Check (4)
    • Finished checking D/A, A40, A30
      • D:ESEP, A30 IP fixed
  • D:ISEP Shielding
    • Cutting steel, staging (broken dumbwaiter), started
  • Ion Chamber removal
  • Deb DCCT motorized stand
    • Will move GAPMON?
  • D:EKIK ceramic tube
  • AP2 Water Hose Replacement
  • Leaking LQA
    • Magnet out of beam line
  • Replaced leaking Deb Sextuple
  • SMA water leak – replace?

Cryo

  • Valve Stem replace
  • Leak Check (2)
  • LN2 transfer line maintenance

Target Station

  • Water Systems (2)
    • DI cartridge replaced
  • Temperature Monitoring (2)
  • Target Rotation
  • Storage
  • SEM Module
  • LVDTs
  • Sample Retrieval
  • Lens motion
  • Air Exhaust Stack
  • 10mm-2 into transformer
    • Prepping
  • 10mm-3 into transformer
    • Prepping
  • Lens replacement
  • Pictures of PMAG

Cooling Systems

  • Deb 3&4 Trans Notch Filters
    • Installed; updating controls
  • Deb Double Notch Filters
    • Installed; updating controls
  • Stacktail 4-8GHz Equalizer
    • Installing
  • TWT balance & trip points
  • Stacktail Filter 3 cable change

Electrical (Pbar & EE Support)

  • Shunt Move
    • Started
  • Shunt Repairs (6)
  • PS Maintenance

SwitchYard

Continuing with maintenance work.

NuMI

  • Continue water sampling through August 20th
  • Drain cleaning completed in Absorber Passageway, Absorber Hall, and Muon Alcoves
  • Updated Fire Protection System Software @ MI65/MINOS Service Buildings
  • Installed 40 High Frequency Cores on Horn Power Supplies to Reduce Noise
  • Subcontractor Dehumidification
    • Installed 4 Support Brackets & 4” Chilled Water Pipes in Shaft
    • Layout 4” chilled water pipes in Upper Utility Labyrinth
    • Started laying out 4” water pipe in Target Hall
    • Spot new chillers in MI65 Service Building Mezzanine
    • Started laying out double wall PVC condensate line for shaft

MiniBooNE

There was no work completed last week.

  • Reverse Horn Polarity
    • Reverse polarity of horn power supply.
    • Test horn.
  • Check Corrosion Tabs at 25m Absorber
    • Unstack 9 blocks from 25m absorber.
    • Retrieve corrosion tabs.
    • Restack blocks.
    • Seal cover.
  • Dry out horn (09Aug07)
  • Check Leakage Current on Ion Pumps
  • Power Supply Maintenance
  • Toroid Work

MTA Beamline Installation

  • Tunnel activities
    • Janitors cleaned up the garage area
    • Contractors started working on the cable tray installation
    • Move an existing turbo cart out of the way
      • Some beam line components have to be installed on that area
  • Shielding under hatch area
    • Parking lot has been roped off since Monday in preparation to receive the crane…well it never came!
    • Iron workers are on site since yesterday
      • And waited all day for the crane which did not show up
        • The delay was due to problems getting the Transportation Permit from the County.
    • Crane is expected to arrive today
      • Activities will begin as soon the mandatory orientation and equipment setup are complete.

Next Week

  • Immediately start the shielding work
    • Estimate 2 days to remove the existing blocks
      • Once this is done, alignment need couple of hours to connect MTA to Linac network
      • Beam line components will then be lowered to the enclosure
  • Tunnel activities continue with contractors installing cable trays
  • Currently the project is 3 days behind schedule. Too early to evaluate the impact of this delay on the overall shutdown schedule.

Collider Detector at Fermilab

  • Flammable gas flow to wire chambers switched to N2
  • Plugs open
  • Preparation proceeding for silicon cooling repair
  • Calorimeter trigger upgrade review
  • Committee signed off on completion
  • Decommissioning of old system will begin next week
  • As found LBQ survey finds changes, meeting next week to discuss repositioning

D-Zero

They have opened up their detector and begun work.

Alignment

Completed Jobs

  • Recycler Task Manager: Oshinowo Field Crew: O’Boyle, Adkins, Teafoe
    • 5527 Optically align Recycler DCCT after installation. Date: 8/13/2007 Complete
  • Tevatron Task Manager: Bocean Field Crew: Bradford, Teafoe, Volk. Crutcher, Adkins, Shaddix
    • 5494 As found low beta quads and separators at B0 Date: 8/06/2007- 8/08/2007 Complete
    • 5495 As found low beta quads and separators at D0. Date: 8/09/2007- 8/14/2007 Complete Task Manager: Bocean Field Crew: Bradford, J Wilson, Watt.
    • Un roll D16-1 quad when D-1 is warmed. (As Found) Date:8/13/2007 Complete
    • 5591 As Founds of the Roman pots at D0 4 of 6 Date:8/13/2007 85% Complete
  • Booster Task Manager: Oshinowo Field Crew: Wilson, Crutcher, Wish.
    • 5534 Align 24 Booster Corrector Magnets after installation. Re-align adjacent cavities. Long 19 Long 17 Date: 8/13/2007 Complete
    • 5534 Align 24 Booster Corrector Magnets (Long 21 Long 23)Date: 8/14/2007 Complete
  • Main Injector Task Manager: Oshinowo Field Crew: O’Boyle, Adkins, Teafoe.
    • 5563 OT. AS found survey of Quadrupoles at 400, 402, 404 and the three Lambertson at MI40. Date: 8/13/2007- 8/15/2007 Complete
    • 5554 307, 308 Base plate survey Date: 8/16/2007 Complete
  • CMS Task Manager: Oshinowo Field Crew: O’Boyle.
    • 5516 Survey the CMS F-pixel Half cylinder +Z2 in the Lab C Date: 8/09/2007 Complete
  • D0 Task Manager: Oshinowo Field Crew: O’Boyle, Adkins, Wish / Wilson, J Wilson
    • Stickmike survey of Muon system prior to opening of detector. This will include both CFs and EFs. Date: 8/06/2007 Complete
    • 5524 VStar or Laser Tracker survey of North/South EF pedestals with respect to the sidewalk monuments. Date: 8/07/2007 Complete
    • 5592 Elevation run in D0 before machine is opened. Date: 8/06/2007 Complete
  • Linac Task Manager: Oshinowo Field Crew: Wilson, Contract Crew.
    • 5468 Stake out the MuCool beamline component locations in the Linac and MuCool Test Area (MTA). Date: 8/08/2007 Complete
  • P-Bar Task Manager: Oshinowo Field Crew: Wilson, O’Boyle, Wyatt.
    • A4Q10 (LQA in Accumulator) has a water leak. An "as found“ Date: 8/09/2007 Complete

Next week’s schedule of work

  • Tevatron Task Manager: Bocean Field Crew: Bradford, Adkins, J Wilson.
    • 5498 Measure rolls on 43 dipoles and quads in Tevatron. These are magnets that cannot be measured with roll fixture. Date: 8/14/2007- 8/27/2007 Field Crew: Crutcher, Wish, Wilson.
    • 5504 As found Pbar kickers in transfer hall (6 components) Date: 8/17/2007 Field Crew: Crutcher, Wish, Wilson.
    • 5505 As found Proton kickers in transfer hall (6 components) Date: 8/20/2007
    • Task Manager: Kyle Field Crew: Contract crew (Smego, Christopher), Wyatt, Volk Elevation run A-sector Complete Date: 8/14/2007- 8/28/2007
  • Booster Task Manager: Oshinowo Field Crew: Wilson, Crutcher, Wish.
    • 5534 Align 24 Booster Corrector Magnets Date: 8/20/2007
  • Main Injector Task Manager: Oshinowo Field Crew: O’Boyle, Adkins, Teafoe.
    • 5554 4 collimators in the Main Injector at the 301, 303, 307, and 308 locations. Survey will be required to survey the Base Plate. Date: 8/17/2007-8/22/2007
  • Nova Task Manager: Oshinowo Field Crew: Wilson, Crutcher, Wish. Date: 8/23/2007
    • 5476- Stake out new Recycler extraction line (MI30) for the Nova Project
    • 5475- Stake out new Recycler injection line (MI10) for the Nova Project
    • 5477 Stake out new Recycler RR30 quad positions for the Nova Project
  • MTA Task Manager: Oshinowo Field Crew: Wilson, Crutcher, Wish.
    • Control Network Continue, including walls and floor centerline layout. Date: 8/21/2007
  • CDF Task Manager: Bocean Not Scheduled
  • ILC Task Manager: Bocean Not Scheduled
  • P150 Task Manager: Oshinowo Field Crew: Wilson, Crutcher, Wish. Late Request
    • 5593 We had to move several magnets and multiwires to make up the ends on the cans for vacuum connections. They are as follows: The multiwire cans at 708 and 710. B2 magnets at the interface of 710 - 711 and the interfaces of 709 - 710, the serial #'s are BDM 883-1 BDM 835-1 BDM 879-1 and BDM 397-1. Date: 8/23/2007

Alignment and Metrology Home page

  • http://ppd.fnal.gov/align/align.html

Mechanical

Janitorial work in the tunnels continues.

Cryogenic System

Tevatron at LN2 temperature except C4 and D1 at room temperature.

  • E1 warm-up to room temperature started.
  • Tunnel magnet cryostat vacuums at A25 and D35-36 regions are of concern.
  • Pbar Cryogenic Transfer Line AP30-AP10/60 vacuum leak checked fine.
  • CDF back on their BA-1 Mycom compressor (failed unit was replaced) and thus isolated from Tevatron Cryogenic System.

Cryo Operations

  • 3 FRIG Mycom compressor (plus DZero & CDF Cryo) running with helium precooled to LN2 temperature flowing through magnet strings.
  • CHL helium inventory management system operating (Tank Farm, Helium Purifier, A0 GHe Recovery) and Tank Farm has been purified; the helium losses are normal.
  • CHL Nitrogen Reliquefier plant off and LN2 supply dependent on vendor deliveries to CHL LN2 Dewar tanks; deliveries are fine.

FRIG Maintenance

  • 7 out of 13 wet expander flywheels have been replaced
  • Wet expander overhauls of cold ends have started
  • CDF BA#1 compressor unit is being replaced (failed last month)
  • Various vacuum, instrumentation, and controls problems being worked on.

CHL Maintenance

  • Coldbox-1:
    • Oil coolers have been replaced
    • Charcoal Adsorber east unit is in high temperature regeneration
  • Nitrogen Reliquefier plant maintenance in progress:
    • Marley Cooling Tower cleanup/overhaul,
    • Rotoflow compressor after cooler heat exchanger tube bundle being replaced
  • Various facility maintenance in progress

Controls

This Week

  • AD Server Updates complete
  • Database Table Update complete
  • MAC & Computer Room UPS Maintenance complete
  • Acc. Network Switch IOS Updates complete (120)
  • Computer Room Prep work for 8/18 Switch replacement
  • MADC II reset updates continue
  • CAMAC Front End SLD Updates continue
  • Continuing routine maintenance
  • JAVA Engine Restarts/Updates
    • Having to reinstall OS on Solaris workstations

Next Week

  • Saturday 8/18
    • Replace EOL Gigabit Network Switch with Cat6506 Chassis
      • No AD general networking (desktops). 2-4hrs.
  • Tuesday 8/21
    • General Network (desktop) Switch IOS Updates
    • VAX Clusters move to MAC Room
      • 4 clusters, move one at a time, down 30 min/cluster
      • Alarms server, Fileshare, Console page access
  • Wednesday 8/22
    • General Network Switch IOS Updates
    • Upgrade Sup720 to Cat6509
      • No Accelerator Division Network. 2-4hrs.
  • Thursday 8/23
    • General Network Switch IOS Updates
  • Saturday 8/25
    • AD Web Server Updates ~2hr.
      • No Time Server, No Elog access, No node reboot capability

Get Requests In Now

  • Hardware you want to use may not be readily available
    • New infrastructure (links, clocks, networking) takes time
  • Work has to get scheduled

FESS

Tevatron Cooling ponds

  • Divers have examined the silt buildup around intake pipes
    • Dredging will begin today
  • A1 pond will have a bypass dug to the next pond
  • This work will affect the cryo in the zero buildings

Electrical

  • Switchyard
    • Many mice nests found, but no live mice
  • Thermography
    • Examined several buildings and found a main buss problem at E1

A/C

  • Work on Tevatron ring dehumidifiers
  • Strike affect minimal

Mechanical

  • Running a fire hose to CHL for CDF ICW bypass
  • Casey’s pond 16” valve work will require a site-wide ICW shutdown

Safety

  • Safety System Status
    • Upgrade M02 - M05 Enclosure Interlocks 12%
    • Upgrade Main Injector PLC Code and replace PLC Batteries 35%
    • Replace Safety System DEC talk with ATT Natural Voices 15%
    • Linux kernal 2.4 to 2.6 upgrade to Safety Data Acquisition System 25%
    • Upgrade Meson Area Oxygen Monitors 0%
    • Install Radiation Monitor Cables for the Muon Test Area 0%
    • Replace Pbar Vault Key Tree Door Latches & add status LEDs 0%
    • Route Key Tree release signals from MCR to Meson 0%
  • A successful shutdown is getting our work done safely

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 08/17/2007    email Fermilab